Calculate the number of grams/serving of sugars in the commercial beverage (Root Beer from the procedures) as measured with the experimental data provided.

Hint: You will need the serving size volume from the can (pictured in procedures) as well as to use your percent concentration of Root Beer found from standard curve in #3 to solve this.

volume=355 mL approx. concentration of root beer=12.5% sucrose
Answer #1

given volume=355ml=0.355l

concentartion 12.5%=0.125 mol/litre

multiply volume with concentarion we get no ofmoles of sucrose:

no of moles=0.355*0.125=0.044375

now multiply no of moles with molar mass of sucrose=

molar mass of sucrose=C12H22O11=342.3 g/mol

no of grams=342.3*0.044375=15.189